OpenStudy (anonymous):

xabs(x-6) = 45x+7. Solve for x.

OpenStudy (freckles):

\[x |x-6|=45x+7\] right?

OpenStudy (anonymous):

yeah

OpenStudy (freckles):

\[|x-6|=\frac{45x+7}{x}\]

OpenStudy (freckles):

is that cool with you I divided by x

OpenStudy (anonymous):

y

OpenStudy (anonymous):

*yup

OpenStudy (freckles):

now we have two equations to consider

OpenStudy (freckles):

since |x-6|=x-6 if x>6 and |x-6|=-(x-6) if x<6

OpenStudy (freckles):

that is we have the following two equations to solve: \[x-6=\frac{45x+7}{x} \\ -(x-6)=\frac{45x+7}{x}\]

OpenStudy (freckles):

the main reason I divided both sides by x initially is that |x-6| can only result in a positive output so that means (45x+7)/x should be positive

OpenStudy (freckles):

now what I'm going to up doing with these two equations is multiply the x again on both sides

OpenStudy (freckles):

\[x(x-6)=45x+7 \text{ or } -x(x-6)=45x+7 \]

OpenStudy (freckles):

now when we solve both of these x's we should verify them by plugging them into the initial equation

OpenStudy (freckles):

both of these equations are quadratic equations

OpenStudy (freckles):

they may not look it to you right now but if you use the distributive property a little bit you will see it

OpenStudy (freckles):

yuck these result in some ugly some answers

OpenStudy (anonymous):

the two quadratic equations you get are x^2 -51x -7 and -x^2 -39x +7 right?

OpenStudy (anonymous):

i meant x^2 -51x -7 and -x^2 -39x -7

OpenStudy (freckles):

yes \[x(x-6)=45x+7 \\ x^2-6x=45x+7 \\ x^2-6x-45x-7=0 \\ x^2-51x-7=0 \\ \text{ other equation is } -x(x-6)=45x+7 \\ -x^2+6x=45x+7 \\ -x^2+6x-45x-7=0 \\ -x^2-39x-7=0 \\ x^2+39x+7=0\]

OpenStudy (anonymous):

so what do you do next?

OpenStudy (freckles):

solve both quadratics

OpenStudy (freckles):

I would probably end up using the approximation for the exact solutions (solutions may be extraneous or not at this point) that is why I'm using the approximations to find out if it is or isn't... so after I find all four exacts I will find approximation to each of those exacts and then plug them in to the (45x+7)/x thing to see if they are positive or not

OpenStudy (freckles):

it looks like you should wind up with 3 solutions

OpenStudy (freckles):

if i didn't make a mistake with my calculator

OpenStudy (freckles):

You can tell me how far you have gotten or show me anything you want from this problem and i can tell you if that is what i got

OpenStudy (freckles):

or if you just want to check something

OpenStudy (freckles):

like what approximations for the 4 numbers are you using?

OpenStudy (mathstudent55):

\(~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~x|x-6| = 45x+7\) \(x(x-6) = 45x+7~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~x[-(x-6)] = 45x+7\) \(x^2-6x = 45x+7~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~-x^2 + 6x = 45x+7\) \(x^2-51x -7 = 0~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~-x^2 - 39x -7 = 0\) \(x^2-51x -7 = 0~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~x^2 + 39x +7 = 0\) \(x = \dfrac{51 \pm \sqrt{(-51)^2 - 4(1)(-7)}}{2(1)} ~~~~~~~~~~~~~ x = \dfrac{-39 \pm \sqrt{(39)^2 - 4(1)(7)}}{2(1)}\) \(x = \dfrac{51 \pm \sqrt{2629}}{2} ~~~~~~~~~~~~~ x = \dfrac{-39 \pm \sqrt{1493}}{2}\) Now you need to check each of the four above solutions to see if it works in the original equation.

OpenStudy (freckles):

have to give you points @mathstudent55 for that cute layout